OpenStudy (anonymous):

what is the change of entropy for a heat engine using 500.0 J at 20.0°C ? 0.04 J/K , 0.25 J/K , 0.59 J/K , or 1.71 J/K ? :/

OpenStudy (michele_laino):

it is simple, we have to apply this formula: \[S = \frac{{\Delta Q}}{T}\]

OpenStudy (anonymous):

okay! 500/20 ? = 25? so our solution would be 0.25 J/K ?

OpenStudy (michele_laino):

no, we have to compute the absolute temperature first: T =20+273=293 °K

OpenStudy (anonymous):

ohh okay! so 500/293 = 1.706? so our solution is 1.71 J/K ?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay! thank you!!

OpenStudy (michele_laino):

:)